# Multimodal classification of forest biodiversity potential from 2D orthophotos and 3D airborne laser scanning point clouds

### v2

@@ 2.5 Dataset creation process
    Filtering the HNV forest map (seen in Figure 3) to include only forest areas with scores from 1–3 or 7–10, used for extracting low forest biodiversity potential samples and high forest biodiversity potential samples, respectively.
@@
    Filtering narrow forest areas using morphological opening and exclude forest areas with an area below 0.5 hectares using connected components analysis, thereby following the minimum forest size definition of FAO and UNEP (2020).
@@
    Randomly selecting 130 low forest biodiversity potential areas and 130 high forest biodiversity potential areas from the filtered areas.
@@
Train High Bio. Potential 88 15 5071550715\,50715 507 (34.9%) 176.2 19.9
Low Bio. Potential 95 11 4181141811\,41811 418 (25.7%) 120.2 13.8
@@ 3 Method
In this section, we describe our methodology for assessing the forest biodiversity potential of samples, using various deep learning methods. We explore five distinct methods: 2D orthophoto classification using ResNet (He et al., 2015), 3D ALS point cloud classification using PointVector (Deng et al., 2023) and three multimodal fusion methods that combine the 2D- and 3D modalities. An architectural overview of all methods are shown in Figure 7.
Figure 7: Classification method overview. Architectural overview of the methods used for classifying samples into high- and low forest biodiversity potential. (a) 2D orthophoto classification using ResNet-18, (b) 3D ALS point cloud classification using PointVector-S, (c) Ensemble classification combining prediction probabilities through weighted averaging, (d) MLP Fusion trained on pre-computed feature embeddings from both modalities and (e) MLP Fusion trained end-to-end.
@@
For the classification of forest biodiversity potential of samples using strictly the 2D orthophotos, we employ the ResNet architecture (He et al., 2015). ResNet is a deep convolutional neural network (CNN) architecture known for its ability to train effectively by using residual connections to mitigate the vanishing gradient problem. The architecture is well-suited for image classification tasks where the model must learn complex patterns from large-scale image datasets. A ResNet consists of a combination of convolutional layers, batch normalization, and ReLU activation functions, followed by a fully connected layer and a softmax layer, as visualized in Figure 7 (a). While newer architectures for 2D image classification, such as the Vision Transformer (ViT) (Dosovitskiy et al., 2021) and ConvNeXt (Woo et al., 2023), demonstrate state-of-the-art performance on many benchmarks, ResNet still remains a valid choice, due to its proven performance, computational efficiency, ease of use, and well-understood behavior. We refer to the original work for further details on the architecture.
@@
For the task of classifying the forest biodiversity potential of samples based on 3D ALS point clouds, we employ the PointVector architecture (Deng et al., 2023). PointVector builds upon previous successful point-based networks such as PointNet, PointNet++ (Qi et al., 2017) and PointNeXt (Qian et al., 2022). PointVector is a state-of-the-art point cloud processing model that introduces a novel vector-oriented point set abstraction module to enhance local feature aggregation, allowing for more effective extraction of spatial information from point cloud data. The model has demonstrated SOTA performance on popular 3D point cloud benchmarks such as S3DIS (Armeni et al., 2016) and ScanObjectNN (Uy et al., 2019), while using significantly fewer parameters than comparable models (Hansen et al., 2024). An architectual overview of PointVector is shown in Figure 7 (b). PointVector extracts features directly from raw 3D ALS point clouds without manual feature engineering. However, the architecture requires a fixed number of input points (e.g., 16 3841638416\,38416 384). Since the ALS point clouds in the BioVista dataset often exceed 30 0003000030\,00030 000 points, down-sampling is necessary, presenting a limitation as potentially relevant spatial information from discarded points is neglected. Further details on the PointVector architecture and its inner workings are beyond the scope of this paper and can be found in the original work.
@@ 3.3.1 Confidence-based ensembling
Our first fusion method, visualized in Figure 7 (c), combines the predictions from the individual models already trained: the ResNet model (for 2D orthophotos) and the PointVector model (for 3D point clouds). This technique is a form of confidence-based ensembling/stacking (Wolpert, 1992).
The process works as follows: For each input sample (which consists of a paired 2D orthophoto and 3D ALS point cloud for a specific forest sample), we first get the predictions from both models independently. Each model outputs a probability score indicating its confidence that the input sample belongs to either the high- or low forest biodiversity potential class.
@@
$P(\text{class}\,|\,\text{input})=w_{\text{2D}}\cdot P_{\text{2D}}(\text{class}\,|\,\text{input})+w_{\text{3D}}\cdot P_{\text{3D}}(\text{class}\,|\,\text{input})\enspace,$
@@ 3.3.2 MLP fusion
MLP Fusion (pre-computed): In this approach, illustrated in Figure 7 (d), the ResNet-18 and PointVector-S backbones act solely as fixed feature extractors. We first process the entire BioVista training, validation and test datasets through the frozen (non-trainable) backbones to generate and store the 512-dimensional feature embeddings for every sample. Subsequently, only the MLP classifier is trained, validated, and evaluated using these pre-computed, static feature vectors. The backbones themselves are not active or updated during this MLP training phase.
@@ 4 Experiments
To evaluate the effectiveness of our proposed methods for classifying forest biodiversity potential using remote sensing data, we conducted a series of experiments using the BioVista dataset. Our experiments aimed to compare the performance of 2D orthophoto-based classification, 3D ALS point cloud-based classification, and fusion approaches that combine both data modalities.
@@ 4.1 Evaluation metrics
Overall accuracy (OAcc) is the most straightforward measure of a classifier’s performance. It is defined as the ratio of correctly classified samples, $N_{\text{correct}}$, to the total number of samples, $N_{\text{total}}$:
$\text{OAcc}=N_{\text{correct}}/N_{\text{total}}$
While OAcc provides a general sense of the model’s performance, it can be misleading in cases of class imbalance. Mean accuracy (MAcc) addresses the potential bias in overall accuracy due to class imbalance and is calculated as the average of the accuracies for each class, thereby, weighing the significance of each class equally important. For BioVista, we calculated the accuracy for classifying low forest biodiversity potential samples, $\text{Acc}_{\text{low}}$ and accuracy for classifying high forest biodiversity potential samples, $\text{Acc}_{\text{high}}$ and define the MAcc as follows:
@@
For the classification of forest biodiversity potential using 2D orthophotos, we employed the ResNet-18 architecture (He et al., 2015). Deeper variants such as ResNet-34 and ResNet-50 were explored in preliminary experiments, but they did not yield significant performance improvements, leading us to select the more computationally efficient ResNet-18.
We set the maximum training epochs to 20, as initial training experiments showed a validation performance plateau after approx. 10–20 epochs of training. We used cross-entropy loss and employed an AdamW optimizer and a cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017, 2016).
@@
The input images have a resolution of 240×240 pixels, maintaining the original 12.5 cm ground sampling distance of the orthophotos. As we employed the sample plot size and shape standards of the Danish NFI (Alban et al., 2019), pixels outside the 30-meter diameter circle are set to zero for all color channels.
To enhance model robustness and prevent over-fitting, we evaluated several data augmentation techniques during training, with inspiration from (Li et al., 2023) and (Tolan et al., 2024):
@@
    Random Gaussian blur with sigma values (0.1–2.0) and kernel size of 5$\times$5 pixels, which could help improve model robustness to image quality differences.
@@
When averaged over seven runs with the best identified hyperparameter settings and trained and evaluated on the RGB-color channels of the orthophotos, ResNet-18 achieved an overall accuracy of $75.5\%$ on the BioVista test set and Table 6. The model performed notably better on the high forest biodiversity potential class ($86.4\%$) compared to the low forest biodiversity potential class ($64.9\%$), indicating a significant class imbalance in the predictive capabilities of the model, which we address in Section 5.
Figure 8: Color channel performance comparison. This figure shows the overall accuracy of a ResNet-18 model on the BioVista test dataset, evaluated using various color channel combinations from orthophotos. The Near-infrared channel performed best among single channels. Additionally, three-channel combinations that included near-infrared (N) achieved higher accuracy than the standard red, green and blue (RGB) combination.
@@ 4.2.1 Near-infrared for forest biodiversity assessment
In addition to conducting experiments with 2D orthophotos using the RGB color channels, we further evaluated ResNet-18 trained on orthophotos with various combination of color channels including the near-infrared channel, as show in Figure 8. When comparing models trained on a only a single color channel, R, G, B or NIR, we see the best performance using the NIR-channel and the worst performance using the red channel.
We observed a clear pattern – models trained on orthophotos with color channels which includes NIR outperforms models trained on color channels excluding the NIR channel. All of the results seen in Figure 8 are averages of at least 7 runs where each run is trained for 20 epochs, using the same hyperparameter settings as described in Section 4.2. We saw the best performance among the evaluated combinations to be models trained on NIR, green and blue channels (NGB), with an overall accuracy of $76.7\%$, as seen in Table 6. For the remainder of this work, 2D ResNet-18 will refer to instances trained on NGB.
@@
For the classification of forest biodiversity potential using 3D ALS point cloud, we employed the PointVector architecture. We used the x, y, z coordinates from the ALS point clouds. Additionally, we added the relative height (h) as an input feature, similar to (Thomas et al., 2019), derived by subtracting all z coordinates of a point cloud with its min z-value.
We trained PointVector for 20 epochs, using cross-entropy loss and the AdamW optimizer and cosine annealing learning rate scheduler, similar to the original PointVector paper Deng et al. (2023).
@@
We carefully evaluated combinations of the hyperparameter settings seen from Table 4 which also displays the final selected values. The batch size and weighted loss settings had a very minor impact on validation performance, while learning rate, query ball radius, and query ball radius scaling had a large impact on the validation performance, with overall accuracy on the validation set ranging from 71.6% with the worst settings to 76.1% with the best settings. For the query ball radius, we initially set it to 0.7 using the method of (Jensen et al., 2023) and evaluated nearby values, resulting in the final selection of 0.65. The performance of experiments in which we sampled 24.576 points from each point cloud actually performed slightly better than for 16.384. However, we chose to go with sampling 16.384 points, as we deemed the trade-off between performance gains and the additional required computational cost not worthwhile.
We evaluated the data augmentation techniques employed by the original PointVector and PointNeXt paper (Qian et al., 2022), which include random rotation around the z-axis (up to 360 degrees), random scaling, and random point jittering. Surprisingly, we found that none of the augmentations had a positive influence on the performance, resulting in the fact that they were skipped for the final evaluation.
Using the hyperparameter settings found in Table 4, the PointVector-S model achieves an overall accuracy of $75.8\%\pm 0.4$ on the BioVista test dataset, when averaged over 7 runs, as seen in Figure 10 and Table 6, which was the lowest performance achieved by the tested methods.
@@
To classify forest biodiversity potential using both 2D orthophotos (NGB channels) and 3D ALS point clouds, we evaluated the three multimodal approaches outlined in Section 3: confidence-based ensembling, MLP fusion with pre-computed features, and end-to-end training.
Confidence-based ensembling: For the confidence-based ensembling approach (Figure 7 (c)), we combined the confidence outputs from seven distinct instances of trained ResNet-18 models and seven distinct instances of trained PointVector-S models. We used weighted averaging, where the weights ($w_{\text{2D}}$ and $w_{\text{3D}}$) were optimized based on performance on the BioVista validation set, as illustrated in Figure 9. Generally, this optimization resulted in a slight weighting preference towards the 3D PointVector predictions, with typical weight ratios ranging from 45:55 to 40:60 (ResNet:PointVector). The weight ratio is surprising, as PointVector evaluated on 3D ALS point clouds achieved worse performance than ResNet evaluated on 2D orthophotos. As this method leverages the outputs of already trained models, no additional training phase was required beyond determining the optimal weights. This ensemble approach achieved an overall accuracy of $80.5\%\pm 2.0$ and a mean accuracy of $80.6\%\pm 1.9$ on the test set. It demonstrated the strongest performance on the high forest biodiversity potential class, achieving $89.5\%\pm 2.7$ accuracy (Table 6).
@@
MLP Fusion (pre-computed): For the MLP Fusion approach using pre-computed features (Figure 7 (d)), we first extracted 512-dimensional feature vectors from the frozen backbones of the trained ResNet-18 and PointVector-S models for all samples in the training, validation, and test sets. We then conducted extensive hyperparameter tuning for the MLP classifier using these static features. The evaluated hyperparameters, their tested ranges, and the final selected values that yielded the best performance on the validation set are summarized in Table 5.
Table 5: Hyperparameter tuning of the MLP Fusion (pre-computed) classifier. Overview of the hyperparameters evaluated, the ranges explored, and the final selected values yielding the best validation performance.
@@
Training with the optimized settings detailed in Table 5 was conducted for 20 epochs using the AdamW optimizer. The pre-computed MLP fusion model yielded the best overall performance among all evaluated methods, achieving an overall accuracy of $81.4\%\pm 0.5$ and a mean accuracy of $81.3\%\pm 0.5$. Notably, this approach produced a more balanced performance between classes compared to the single-modality models and the ensemble method, with $85.8\%\pm 4.5$ accuracy on the high potential class and $76.9\%\pm 3.8$ accuracy on the low potential class, as detailed in Table 6 and visualized regionally in Figure 10.
MLP Fusion (end-to-end): Finally, we evaluated the end-to-end MLP Fusion approach (Figure 7 (e)), where both the MLP classifier and the ResNet-18 and PointVector-S backbones were trained simultaneously. For the MLP classifier part, we utilized the same optimal hyperparameters identified for the pre-computed approach (Table 5). The backbones themselves were fine-tuned during this process. We experimented with different learning rates for the backbones, testing $10^{-4}$ (matching the MLP head), $10^{-5}$, and $10^{-6}$, finding that a lower learning rate of $10^{-5}$ for the backbones yielded the best results. Due to memory constraints associated with simultaneous training of all components, we employed gradient accumulation. Gradients were accumulated over 8 steps using a base batch size of 8 to achieve an effective batch size of 64. As shown in Table 6, the end-to-end model achieved an overall accuracy of $80.4\%\pm 0.2$ and mean accuracy of $80.5\%\pm 0.2$. While its overall performance was slightly below the pre-computed MLP fusion, it achieved the highest accuracy on the low forest biodiversity potential class among all methods ($77.3\%\pm 2.1$). Furthermore, the very low standard deviations across the seven runs (e.g., ±$\pm 0.2\%$ for OAcc and MAcc) indicate that this end-to-end training regime resulted in the most stable and consistent model performance. The performance of the end-to-end trained MLP Fusion model (Figure 7 (e)) is also included in Table 6 for comparison and regionally in Figure 10.
Table 6: Overall model performance. The performance metrics of our classification models evaluated on the BioVista test set, including overall accuracy (OAcc), mean accuracy (MAcc), accuracy for the high forest biodiversity potential class ($\text{Acc}_{\text{high}}$), and accuracy for the low forest biodiversity potential class ($\text{Acc}_{\text{low}}$). All metrics are reported as means with associated standard deviation across seven runs with the same hyperparameter settings, trained on the entire BioVista training set.
@@
2D + 3D Ensembling 80.5% ± 2.0 80.6% ± 1.9 89.5% ± 2.7 71.8% ± 6.0
2D + 3D MLP Fusion (pre-computed) 81.4% ± 0.5 81.3% ± 0.5 85.8% ± 4.5 76.9% ± 3.8
@@
4.5 2D-3D multimodal fusion benefits
From Table 6, we see that the 2D + 3D MLP fusion model significantly outperformed the single modality 2D ResNet and 3D PointVector models. The question arises whether the performance gain was not merely due to ensembling two models (e.g., ResNet-8 and PointVector). We verified that the improvement stems indeed from the fact that the 2D orthophoto and 3D ALS point cloud data modalities carry distinct and complementing features. We did this by comparing the performance of MLP models, with settings as described in Section 4.4, trained on outputs from multiple combinations of two ResNet models (2D + 2D) and two PointVector models (3D + 3D) to the performance of the 2D + 3D MLP fusion model. We used three distinct instances of ResNet models and three distinct instances of PointVector models, trained with hyperparameter settings as described in Section 4.2 and 4.3 respectively, which we refer to ResNet 1, ResNet 2 and ResNet 3 and PointVector 1, PointVector 2 and PointVector 3 in Table 7. The ResNet models were trained on orthophotos using the NIR, G and B color channels. Then we compared the performance of standalone models against fused models. We calculated the percentage point gain of the fused model version compared to the average performance of the standalone models and to the best performing standalone model, referred to as gain (avg) and gain (max) in Table 7.

### v3

@@ 2.5 Dataset creation process
    Partitioning the HNV forest map (seen in Figure 3) into areas with scores from 1–3, 7–10, representing low and high forest biodiversity potential samples, respectively.
@@
    Filtering narrow forest areas using morphological opening and exclude forest areas with an area below 0.5 hectares using connected components analysis, thereby following the minimum forest size definition of fao-state-of-worlds-forests-2020.
@@
    Randomly selecting forest areas for each biodiversity class from the filtered areas to avoid class-imbalance, whereby we chose 130 areas per class.
@@
Train High Bio. Potential 88 15 50715\,507 (34.9%) 176.2 19.9
Low Bio. Potential 95 11 41811\,418 (25.7%) 120.2 13.8
@@ 3 Method
In this section, we describe our methodology for assessing the forest biodiversity potential of samples, using various deep learning methods. We explore five distinct methods: 2D orthophoto classification using ResNet (kaiming-resnet), 3D ALS point cloud classification using PointVector (deng-pointvector) and three multimodal fusion methods that combine the 2D- and 3D modalities. An architectural overview of all methods are shown in Figure 7.
Figure 7: Classification method overview. Architectural overview of the methods used for classifying samples into high- and low forest biodiversity potential. (a) 2D orthophoto classification using ResNet-18, (b) 3D ALS point cloud classification using PointVector-S, (c) Ensemble classification combining prediction probabilities through weighted averaging, (d) MLP Fusion trained with frozen backbones and (e) MLP Fusion trained end-to-end.
@@
For the classification of forest biodiversity potential of samples using strictly the 2D orthophotos, we employ the ResNet architecture (kaiming-resnet). ResNet is a deep convolutional neural network (CNN) architecture known for its ability to train effectively by using residual connections to mitigate the vanishing gradient problem. The architecture is well-suited for image classification tasks where the model must learn complex patterns from large-scale image datasets. A ResNet consists of a combination of convolutional layers, batch normalization, and ReLU activation functions, followed by a fully connected layer and a softmax layer, as visualized in Figure 7 (a). While newer architectures for 2D image classification, such as the Vision Transformer (ViT) (dosovitskiy-2021-vistion-transformer) and ConvNeXt (Woo-2023-ConvNeXtV2), demonstrate state-of-the-art performance on many benchmarks, ResNet still remains a valid choice, due to its proven performance, computational efficiency, ease of use, and well-understood behavior. We refer to the original work for further details on the architecture.
@@
For the task of classifying the forest biodiversity potential of samples based on 3D ALS point clouds, we employ the PointVector architecture (deng-pointvector). PointVector builds upon previous successful point-based networks such as PointNet, PointNet++ (qi-pointnet++) and PointNeXt (guocheng-pointnext). PointVector is a state-of-the-art point cloud processing model that introduces a novel vector-oriented point set abstraction module to enhance local feature aggregation, allowing for more effective extraction of spatial information from point cloud data. The model has demonstrated state of the art performance on popular 3D point cloud benchmarks such as S3DIS (armeni-s3dis) and ScanObjectNN (uy-scanobjectnn), while using significantly fewer parameters than comparable models (opentrench3d). An architectual overview of PointVector is shown in Figure 7 (b). PointVector extracts features directly from raw 3D ALS point clouds without manual feature engineering. However, the architecture requires a fixed number of input points (e.g., 16 38416\,384). Since the ALS point clouds in the BioVista dataset often exceed 30 00030\,000 points, down-sampling is necessary, presenting a limitation as potentially relevant spatial information from discarded points is neglected. Further details on the PointVector architecture and its inner workings are beyond the scope of this paper and can be found in the original work.
@@ 3.3.1 Confidence-based ensembling
Our first fusion method, visualized in Figure 7 (c), combines the predictions from the individual models already trained: the ResNet model (for 2D orthophotos) and the PointVector model (for 3D point clouds). This technique is a form of confidence-based ensembling/stacking (wolpert1992stacked).
The process works as follows: For each input sample (which consists of a paired 2D orthophoto and 3D ALS point cloud for a specific forest sample), we first get the predictions from both models independently. Each model outputs a probability score indicating its confidence that the input sample belongs to a particular forest biodiversity potential class.
@@
$P(\text{class}\,|\,\text{input})=w_{\text{2D}}\cdot P_{\text{2D}}(\text{class}\,|\,\text{input})+w_{\text{3D}}\cdot P_{\text{3D}}(\text{class}\,|\,\text{input})\kern 5.0pt,$
@@ 3.3.2 MLP fusion
MLP Fusion (frozen backbones): In this approach, illustrated in Figure 7 (d), the ResNet-18 and PointVector-S backbones act solely as frozen (non-trainable) high-dimensional feature extractors. During training and evaluation, input samples are passed through these backbones to generate 512-dimensional feature embeddings, which are then fed into the MLP classifier. Only the MLP parameters are updated, while the backbone weights remain fixed throughout the process.
@@ 4 Experiments
To evaluate the effectiveness of our proposed methods for classifying forest biodiversity potential using remote sensing data, we conducted a series of experiments using the BioVista dataset. Our experiments aimed to compare the performance of 2D orthophoto-based classification, 3D ALS point cloud-based classification, and fusion approaches that combine both data modalities. The primary focus of our analysis is the binary classification task distinguishing areas of high and low forest biodiversity potential. For completeness, we also performed a three-class experiment incorporating a medium-potential category (HNV scores 4–6) to examine the increased difficulty of distinguishing among high, medium, and low biodiversity levels. This additional experiment is presented only in this section, as we regard the binary setup as the principal and most meaningful configuration. Accordingly, we generated 26 83426\,834 medium-potential samples (train: 16 08016\,080, validation: $5380$, test: $5374$) using the procedure outlined in Section 2.5.
@@ 4.1 Evaluation metrics
Overall accuracy (OAcc) is the most straightforward measure of a classifier’s performance. It is defined as $\text{OAcc}=N_{\text{correct}}/N_{\text{total}}$, the ratio of correctly classified samples, $N_{\text{correct}}$, to the total number of samples, $N_{\text{total}}$. While OAcc provides a general sense of the model’s performance, it can be misleading in cases of class imbalance. Mean accuracy (MAcc) addresses the potential bias in overall accuracy due to class imbalance and is calculated as the average of the accuracies for each class, thereby, weighing the significance of each class equally important. In the binary classification scenario, we calculated the accuracy for classifying low forest biodiversity potential samples, $\text{Acc}_{\text{low}}$ and accuracy for classifying high forest biodiversity potential samples, $\text{Acc}_{\text{high}}$ and define the MAcc as follows:
@@
For the classification of forest biodiversity potential using 2D orthophotos, we employed the ResNet-18 architecture (kaiming-resnet). Deeper variants such as ResNet-34 and ResNet-50 were explored in preliminary experiments, but they did not yield significant performance improvements, leading us to select the more computationally efficient ResNet-18.
We set the maximum training epochs to 20, as initial training experiments showed a validation performance plateau after approx. 10–20 epochs of training. We used cross-entropy loss and employed an AdamW optimizer and a cosine annealing learning rate scheduler (2017-loshchilov-adamw; 2016-loshchilov-cosine-annealing-lr).
@@
The input images have a resolution of 240×240 pixels, maintaining the original 12.5 cm ground sampling distance of the orthophotos. As we employed the sample plot size and shape standards of the Danish NFI (alban-nfi-field-instructions), pixels outside the 30-meter diameter circle are set to zero for all color channels.
To enhance model robustness and prevent over-fitting, we evaluated several data augmentation techniques during training, with inspiration from (li2023treeheightprediction) and (tolan2024canopy-height-maps):
@@
    Random Gaussian blur with sigma values (0.1–2.0) and kernel size of $\times5$ pixels, which could help improve model robustness to image quality differences.
@@
When averaged over seven runs with the best identified hyperparameter settings and trained and evaluated on the RGB-color channels of the orthophotos, ResNet-18 achieved an overall accuracy of $77.6\%$ on the BioVista test set on the binary classification tasktable updated. The model performed notably better on the high forest biodiversity potential class ($88.0\%$) compared to the low forest biodiversity potential class ($67.5\%$), indicating a significant class imbalance in the predictive capabilities of the model, which we address in Section 5.
As expected, this task proved more challenging: the confusion matrix computed over seven trials in Table 7a shows that while the high-potential class can still be identified reliably, the low-potential class is frequently confused with the medium-potential class.
@@ 4.2.1 Near-infrared for forest biodiversity assessment
In addition to conducting experiments with 2D orthophotos using the RGB color channels, we further evaluated ResNet-18 trained on orthophotos using the RGB channels plus the near-infrared (NIR) channel. Including the NIR channel consistently improved performance. On the binary task, ResNet-18 using RGB and NIR achieved an overall accuracy of $77.6\%$, see Table 6.
@@
For the classification of forest biodiversity potential using 3D ALS point cloud, we employed the PointVector architecture. We used the x, y, z coordinates from the ALS point clouds. Additionally, we added the relative height (h) as an input feature, similar to (2019-thomas-hugues-kpconv), derived by subtracting all z coordinates of a point cloud with its min z-value.
We trained PointVector for 20 epochs, using cross-entropy loss and the AdamW optimizer and cosine annealing learning rate scheduler, similar to the original PointVector paper deng-pointvector.
@@
We carefully evaluated combinations of the hyperparameter settings seen from Table 4 which also displays the final selected values. The batch size and weighted loss settings had a very minor impact on validation performance, while learning rate, query ball radius, and query ball radius scaling had a large impact on the validation performance, with overall accuracy on the validation set ranging from 71.6% with the worst settings to 76.1% with the best settings. For the query ball radius, we initially set it to 0.7 using the method by jensen-data-driven-tuning-2023 and evaluated nearby values, resulting in the final selection of 0.65. The performance of experiments in which we sampled 24.576 points from each point cloud actually performed slightly better than for 16.384. However, we chose to go with sampling 16.384 points, as we deemed the trade-off between performance gains and the additional required computational cost not worthwhile.
We evaluated the data augmentation techniques employed by the original PointVector and PointNeXt paper (guocheng-pointnext), which include random rotation around the z-axis (up to 360 degrees), random scaling, and random point jittering. Surprisingly, we found that none of the augmentations had a positive influence on the performance, resulting in the fact that they were skipped for the final evaluation.
On the binary task, using the hyperparameter settings from Table 4, the PointVector-S model achieves an overall accuracy of $75.8\%\pm 0.4$ on the BioVista test dataset, when averaged over 7 runs, as seen in Figure 9 and Table 6, which was the lowest performance achieved by the tested methods. When distinguishing all three forest biodiversity potential classes, the performance dropped significantly, see the confusion matrix in Table 7 b). Both low and high potential areas are often confused with the medium potential class.
@@
To classify forest biodiversity potential using both 2D orthophotos and 3D ALS point clouds, we evaluated the three multimodal approaches outlined in Section 3: confidence-based ensembling, MLP fusion with frozen backbones, and end-to-end training.
Confidence-based ensembling: For the confidence-based ensembling approach (Figure 7 (c)), we combined the confidence outputs from seven distinct instances of trained ResNet-18 models and seven distinct instances of trained PointVector-S models. We used weighted averaging, where the weights ($w_{\text{2D}}$ and $w_{\text{3D}}$) were optimized based on performance on the BioVista validation set, as illustrated in Figure 8. Generally, this optimization resulted in a slight weighting preference towards the 3D PointVector predictions, with typical weight ratios ranging from 45:55 to 40:60 (ResNet:PointVector). The weight ratio is surprising, as PointVector evaluated on 3D ALS point clouds achieved worse performance than ResNet evaluated on 2D orthophotos. As this method leverages the outputs of already trained models, no additional training phase was required beyond determining the optimal weights. This ensemble approach achieved an overall accuracy of $79.9\%\pm 1.1$ and a mean accuracy of $80.1\%\pm 1.1$ on the test set.
@@
MLP Fusion (frozen backbones): For the MLP Fusion approach using frozen backbones (Figure 7(d)), input samples are passed through the non-trainable ResNet-18 and PointVector-S backbones during training and evaluation to generate 512-dimensional feature embeddings, which are then fed into the MLP classifier. Only the MLP parameters are updated while the backbone weights remain fixed. We performed extensive hyperparameter tuning for the MLP using these dynamically generated features. The evaluated hyperparameters, their tested ranges, and the final selected values that yielded the best performance on the validation set are summarized in Table 5.
Table 5: Hyperparameter tuning of the MLP Fusion (frozen backbones) classifier. Overview of the hyperparameters evaluated, the ranges explored, and the final selected values yielding the best validation performance.
@@
Training with the optimized settings detailed in Table 5 was conducted for 20 epochs using the AdamW optimizer. The frozen backbones MLP fusion model yielded the best overall performance among all evaluated methods, achieving an overall accuracy of $81.5\%\pm 1.0$ and a mean accuracy of $81.7\%\pm 0.9$. Notably, this approach produced a more balanced performance between classes compared to the single-modality models and the ensemble method, with $91.3\%\pm 1.0$ accuracy on the high potential class and $72.1\%\pm 2.6$ accuracy on the low potential class, as detailed in Table 6 and visualized regionally in Figure 9.
MLP Fusion (end-to-end): Finally, we evaluated the end-to-end MLP Fusion approach (Figure 7 (e)), where both the MLP classifier and the ResNet-18 and PointVector-S backbones were trained simultaneously. For the MLP classifier part, we utilized the same optimal hyperparameters identified for the frozen backbones approach (Table 5). The backbones themselves were fine-tuned during this process. We experimented with different learning rates for the backbones, testing $10^{-4}$ (matching the MLP head), $10^{-5}$, and $10^{-6}$, finding that a lower learning rate of $10^{-5}$ for the backbones yielded the best results. Due to memory constraints associated with simultaneous training of all components, we employed gradient accumulation. Gradients were accumulated over 8 steps using a base batch size of 8 to achieve an effective batch size of 64. As shown in Table 6, the end-to-end model achieved an overall accuracy of $82.0\%\pm 0.7$ and mean accuracy of $82.2\%\pm 0.7$. The end-to-end model performed overall best and showed the low standard deviation. The low forest biodiversity potential is $91.1\%\pm 0.7$ and high forest biodiversity potential of $73.2\%\pm 1.9$ — nearly matching the frozen-backbone model in the latter case. Hence, it led in all categories except the high-forest score, where results were similar. The performance of the end-to-end trained MLP Fusion model (Figure 7 (e)) is also included in Table 6 for comparison and regionally in Figure 9.
Table 6: Overall model performance on the binary classification task. The performance metrics of our classification models evaluated on the BioVista test set, including overall accuracy (OAcc), mean accuracy (MAcc), accuracy for the high forest biodiversity potential class ($\text{Acc}_{\text{high}}$), and accuracy for the low forest biodiversity potential class ($\text{Acc}_{\text{low}}$). All metrics are reported as means with associated standard deviation across seven model runs with the same hyperparameter settings, which are all trained on the entire BioVista training set.
@@
2D + 3D Ensembling 79.9% ± 1.1 80.1 % ± 1.1 89.6% ± 2.0 70.6% ± 2.1
2D + 3D MLP Fusion (frozen backbones) 81.5% ± 1.0 81.7% ± 0.9 91.3% ± 1.0 72.1% ± 2.6
2D + 3D MLP Fusion (end-to-end) 82.0% ± 0.7 82.2% ± 0.7 91.1% ± 0.7 73.2% ± 1.9
Table 7: Row-normalized confusion matrices of a) ResNet-18, b) PointVector, c) PointVector and ResNet-18 combined with frozen backbones, and d) PointVector and ResNet-18 combined trained end-to-end for three classes and averaged over seven trials.
predicted
high low med.
[…]
a) ResNet-18
predicted
high low med.
[…]
b) PointVector
predicted
high low med.
[…]
c) Fusion (frozen)
predicted
high low med.
[…]
d) Fusion (end-to-end)
@@
4.5 2D-3D multimodal benefits
From Table 6, we see that the 2D + 3D model significantly outperformed the single modality 2D ResNet and 3D PointVector models when considering the binary classification task. As can be seen in Table 7, the combined models also performed better in the three-category setting.
The question arises whether the performance gain was not merely due to ensembling two models (e.g., ResNet-18 and PointVector). We verified that the improvement stems indeed from the fact that the 2D orthophoto and 3D ALS point cloud data modalities carry distinct and complementing features. Concretely, we did this by comparing the performance of ensembling models of two ResNet models (2D + 2D) , two PointVector models (3D + 3D) to the performance of the 2D + 3D model. We used three distinct instances of ResNet models and three distinct instances of PointVector models, all trained with exactly the same hyperparameter settings, which are shown in Table 3 and Table 4 respectively, which we refer to ResNet 1, ResNet 2 and ResNet 3 and PointVector 1, PointVector 2 and PointVector 3 in Table 8updated table. The only difference between the model instances are the random seeds which the model weights are initialized with when the model training starts. The ResNet models were trained on orthophotos using the RGB and NIR channels. Then we compared the performance of standalone models against ensembled models. We calculated the percentage point gain of the ensembled model version compared to the average performance of the standalone models and to the best performing standalone model, referred to as gain (avg) and gain (max) in Table 8.
Table 8: Multimodal ensemble performance comparison. Analysis of model performance between different modality pairs (2D-2D, 3D-3D, and 2D-3D), showing overall accuracy for individual models and fused versions along with performance gains relative to average and maximum single model accuracies.